if roots of a equation x^2-(a+3)x+3a-1=0 are integarl then the value of a is 

Grade:12th pass

1 Answers

jagdish singh singh
7 years ago
\hspace{-0.7 cm}$ Given $x^2-(a+3)x+3a=1\Rightarrow (x-a)(x-3) = 1=1\cdot 1 = -1 \cdot -1 $\\\\ $\bullet$ If $x-a=1$ and $x-3 = 1\;,$ Then we get $a=3$\\\\ $\bullet$ If $x-a=-1$ and $x-3 = -1\;,$ Then we get $a=3$